### v1

# VaLID: Verification as Late Integration of Detections for LiDAR-Camera Fusion

Abstract
Vehicle object detection is possible using both LiDAR and camera data. Methods using LiDAR generally outperform those using cameras only. The highest accuracy methods utilize both of these modalities through data fusion. In our study, we propose a model-independent late fusion method, VaLID, which validates whether each predicted bounding box is acceptable or not. Our method verifies the higher-performing, yet overly optimistic LiDAR model detections using camera detections that are obtained from either specially trained, general, or open-vocabulary models. VaLID uses a simple multi-layer perceptron trained with a high recall bias to reduce the false predictions made by the LiDAR detector, while still preserving the true ones. Evaluating with multiple combinations of LiDAR and camera detectors on the KITTI dataset, we reduce false positives by an average of 63.9%, thus outperforming the individual detectors on 2D average precision (2DAP). Our approach is model-agnostic and demonstrates state-of-the-art competitive performance even when using generic camera detectors that were not trained specifically for this dataset.
I INTRODUCTION
The development of autonomous vehicles relies heavily on accurate object detection to safely navigate their environment. These systems use multiple sensors, such as cameras and LiDAR, each offering unique strengths and weaknesses. Cameras provide detailed visual information, while LiDAR offers precise depth information crucial for 3D localization. While systems using only cameras or LiDAR data provide valuable information, single-sensor approaches often fail. Cameras struggle in low-light conditions [23, 5], and LiDAR lacks rich high resolution visual data [51, 26, 12, 32].
Multi-modal fusion is often used to ensure robust and reliable detection [36, 37]. These methods combine LiDAR’s depth data with camera visuals to create a more reliable detection system. However, the integration of these modalities often requires advanced fusion techniques with methods designed to match the specific characteristics of the underlying data [40, 10, 50, 1]. The primary challenge is ensuring that the strengths of each modality complement one another to improve overall detection accuracy in varied and dynamic driving conditions.
Sensor fusion in autonomous vehicles can be categorized into early, deep, and late fusion techniques. Early fusion integrates raw data from multiple sensors early in the processing pipeline, allowing for detailed cross-sensor interaction [34, 47]. However, this approach can lead to increased computational costs due to the need for complex preprocessing steps, like semantic segmentation or depth completion [45, 42, 7]. Deep fusion aligns features at a higher abstraction level, balancing information from both sensors but adding complexity to the model design[35, 20]. Late fusion, in contrast, merges detection outputs at the bounding box level after independent processing, offering more flexibility and computational efficiency but with reduced feature interaction [24, 25].
Early and deep fusion methods leverage cross-modality information in their fundamental representations, potentially leading to higher accuracy at the cost of complexity. Late fusion methods, on the other hand, are simpler to integrate into complex industrial workflows, since they do not require a redesign when advances are made in underlying detection technologies. Pre-trained single-modality detectors can be used and replaced without modification, requiring only detection-level outputs at the time of fusion. An ideal fusion method would allow the flexibility of late-fusion while obtaining the performance of early and deep fusion.
We introduce a new late fusion method called VaLID: Verification as Late Integration of Detections. This method takes raw detections from a LiDAR model and cross-validates them against corresponding camera detections using a simple multi-layer perceptron neural network. Ensuring that only verified LiDAR boxes are kept significantly reduces false positives, improving overall performance.
We evaluate our method on the KITTI dataset with two LiDAR detectors, PV-RCNN[31] and TED-S[41]. To ensure generality, we use three distinct camera models: MonoDETR[48] is a special purpose detector trained specifically on KITTI, YOLO-NAS is a general-purpose model from the YOLO[29] family, and Grounding DINO[16] is an open-vocabulary model.
Detection methods trained for a specific dataset normally outperform general purpose models. Our two LiDAR methods were selected as state-of-the-art methods designed and tested on the KITTI dataset. For camera detectors we included general purpose detectors in our evaluation. These have lower performance than a specialized model when considered as a primary detector, however our method uses camera models for the easier task of verification.
We find that the VaLID method removes an average 63.9% of false positives, consistently improving overall detection average precision. Importantly, improvement is as effective using the general camera models as it is with the specifically trained method. In both cases results are competitive with state of the art.
The contribution of this paper is a model-independent late fusion method that combines LiDAR and camera detections.
Figure 1: We conduct our study on the KITTI dataset. The figure rows show (a) official 2D ground truth (b) detections from the specialized LiDAR model PVRCNN, (c) specialized camera model MonoDETR, and (d) open Vocabulary model GroundingDino. LiDAR generally produces too many false positives, seen most obviously in Column [i]. Camera models can help verify and reject false positives despite imperfections in their own detections.
II Related Work
II-A Detection Using Camera Images
Cameras images are fundamentally a 2D data source, and 2D object detection is a well-studied area with many surveys and general purpose datasets [52, 21, 38, 16, 17, 28, 30, 15, 6]. Higher accuracy is possible in specific domains such as vehicle detection by designing specialized object detection methods which are trained on vehicle specific datasets [8, 33, 3]. Since autonomous driving requires 3D awareness, there have also been attempts to detect 3D bounding boxes from camera images. However, accurately estimating depth without dedicated sensors is challenging, especially for distant objects. Early approaches like Deep3DBox leverage 2D bounding boxes and geometric constraints to infer 3D dimensions but struggle with long-range accuracy [23]. Similarly, MonoDETR improves detection by incorporating depth embeddings into a transformer model [48]. Other methods generate depth maps to assist with detection. For instance, MonoPGC refines pixel-wise depth estimation through cross-attention [44], and Pseudo-LiDAR simulates LiDAR input from depth maps to enhance detection performance [39]. While these techniques improve depth accuracy, they often introduce significant computational overhead, especially for complex or occluded scenes. Multi-task approaches like Deep MANTA predict both 2D and 3D attributes simultaneously, leveraging shared features to enhance efficiency, but they still face challenges in reliable depth estimation [5]. Unfortunately, camera-only methods, while effective in some scenarios, are limited by their inability to produce reliable depth estimation without additional sensors, and 3D bounding box accuracy remains below 2D bounding box accuracy. In our work, we use the higher accuracy 2D detections as one of the inputs to our late fusion method, which integrates data from multiple sensors.
II-B Detection Using LiDAR Point Clouds
LiDAR-based 3D detection has become a fundamental component in autonomous driving due to its ability to capture detailed spatial information. One-stage models, like PointNet and PointNet++, were early efforts that applied neural networks directly to raw point clouds [26], [27]. These methods prioritized real-time processing but often struggled with long-range accuracy. Other approaches, like those using range images, attempt to enhance detection by encoding depth information at the pixel level, yet they still face limitations in handling complex, distant scenes[22],[2]. To address these challenges, two-stage models were developed to refine detection for better accuracy. VoxelNet [51] and SECOND [46] introduced the use of 3D voxels and sparse convolutions, enhancing both feature extraction and processing efficiency. Building on this, PV-RCNN combines voxel-based CNNs with PointNet abstraction to generate high-quality 3D proposals [31], while PointRCNN [32] refines 3D proposals in a two-stage process. Part-A2Net [31] goes a step further by predicting part locations within objects to refine proposals, improving 3D bounding box accuracy. LiDAR-only methods offer robust spatial data but often struggle with complex environments and distant objects. In our late fusion approach, LiDAR detections are combined with camera detections to provide more accurate and reliable 3D object detection.
II-C Detection Using LiDAR-Camera Fusion
Multi-modal fusion techniques for 3D detection are increasingly common, particularly in autonomous driving, where camera and LiDAR sensors complement each other. These techniques are categorized into three main approaches: early, deep, and late fusion.
Figure 2: Precision-Recall curve (PRC) of our baseline single modality models on the KITTI moderate difficulty data set. Notice that both LiDAR models outperform all three of the camera models, and that the specialized camera detector outperforms the two general purpose camera-based object detectors.
Early fusion methods integrate camera data with LiDAR point clouds early in the perception pipeline to enhance object detection. Techniques like PointPainting and MVP enrich LiDAR data by projecting 2D camera segmentation or generating virtual points to address LiDAR sparsity in distant regions [34, 47]. However, image-based segmentation can blur object boundaries, a problem mitigated by FusionPainting using 3D segmentation [45]. Techniques like VirConv create virtual points and apply Stochastic Voxel Discard to filter noise, though this adds computational overhead [42, 7].
Deep fusion methods operate at a higher level of abstraction, aligning features from both modalities. Models like PointAugmenting and DVF bridge the 2D-3D gap by projecting point clouds onto image feature maps or reweighting voxel features based on pixel confidence [35, 20]. Advanced models such as EPNet, EPNet++, and DeepFusion use attention mechanisms to dynamically adjust feature weights, reducing inconsistencies [11, 18, 14]. SFD addresses information loss by fusing features directly in the 3D Region of Interest (RoI), while BEVFusion unifies features into a Bird’s Eye View (BEV) representation [43, 19].
Late fusion integrates outputs from each sensor after independent processing. For example, CLOCs and Fast-CLOCs refine detection confidence by ensuring that 2D and 3D bounding boxes overlap accurately, but these methods primarily focus on raw data from each sensor[24, 25]. They use Multi-Layer Perceptrons (MLP) to process detection candidates, but insufficient feature alignment can result in suboptimal training performance. C-CLOCs addresses this limitation by incorporating contrastive learning and aligning features from both 2D and 3D sensors to improve detection accuracy, particularly in challenging environments[49]. Çaldıran and Acarman take a different approach, integrating 2D segmentation with 3D LiDAR to reduce false positives, such as roadblocks and tunnel walls, improving detection accuracy [4]. All of these late-fusion methods for vehicle detection have focused on combining specialized LiDAR and camera vehicle detectors.
Our work proposes a new late fusion method that is competitive with the state of the art. We demonstrate results with both a specialized camera detector trained on the KITTI dataset and with more general vision models that lack specific knowledge of the dataset.
Figure 3: The distribution of false positives and true positives across confidence score bands for one LiDAR and two camera object detection models on the KITTI moderate set. Notice that the LiDAR method has false positives across all confidence bands, and that the camera methods have different and sometimes complementary distributions.
Figure 4: In our method called VaLID, we take LiDAR (L) boxes as the primary detections and use the camera (C) modality to verify whether each L box is an acceptable detection or not. The input vector is given as the bounding box dimensions of width $w$, height $h$, and center $(cx,cy)$, confidence scores $s$, and the measure of overlap $IoU$. We pass this vector through an MLP and output an accept or reject signal as a sigmoid value. The target loss is calculated considering the overlap of the ground truth with the given L box.
III Background
III-A LiDAR and Camera Models for Late-Fusion
Many urban driving object detection methods utilize LiDAR and camera modalities separately, as well as in fusion, for enhanced performance. Most methods, however, use specialized training and fine-tuning tailored to specific datasets, which is both time-consuming and computationally demanding. To reduce this dependency, our study leverages general and open-vocabulary camera models without requiring dataset-specific fine-tuning. We perform our study on the KITTI dataset [8], using PVRCNN [31] and TED-S [41] models as the primary LiDAR models. For fusion, we use MonoDETR [48] as a specialized KITTI model, and YOLO-NAS [29] as a general camera detector that is not trained on KITTI. GroundingDINO [16] is used as an open-vocabulary model that can detect objects that are not even seen during the training. Although not as precise as fine-tuned models, these general models offer cross-domain versatility, providing the flexibility needed for a more robust late-fusion approach.
III-B Model Detections Complement Each Other
LiDAR and camera models have their own strengths and weaknesses. Fig. 1 illustrates how these models, while missing certain detections, often compensate for one another’s limitations. For instance, PVRCNN can identify numerous objects but tends to generate a lot of false positives. Meanwhile, MonoDETR and GroundingDINO may miss some detections covered by PVRCNN, but they also detect objects that PVRCNN overlooks. GroundingDINO, despite producing imperfect bounding boxes and some false positives, detects objects that specialized models may miss.
Fig. 2 shows the overall performance of our baseline LiDAR and camera methods. Notice that both LiDAR methods outperform all three camera methods, and that the specialized camera model outperforms the two general models. Since the LiDAR methods outperform the camera methods, we take them as our primary detectors, and use the camera methods to augment their detection ability.
Unfortunately, the tested LiDAR methods produce many false positives which negatively impacts the overall system performance. Reducing these false positives is crucial for enhancing precision. To better understand the distribution of false positives, Fig. 3 shows false positives and true positives over different confidence bands for several methods. Notice that the overall distributions are different in different methods. The neural network in our fusion model learns to leverage these differences.
IV Design and Implementation
IV-A Method Overview
We use camera model detections and a neural network to verify whether a particular LiDAR bounding box detection is valid. Fig. 4 provides an overview of our method. For each LiDAR detection, we look for a corresponding camera detection. If a matching camera box is found, the camera bounding box, the Intersection over Union (IoU) between the LiDAR and camera boxes, and the confidence scores of both detections are fed into a neural network, along with the LiDAR box. If no matching camera box is found, we input a zero for both the bounding box and the confidence score for the camera model.
Our goal is to determine whether to keep or discard each LiDAR (L) detection. We generate ground truth by calculating the IoU between the LiDAR box and the official KITTI ground truth, and follow the datasets standard by setting an IoU threshold of 0.7. Detections above this threshold are marked as valid (positive class), while those below are ignored. The inputs given would be:
$I=\{\text{Bbox}_{L},\text{Bbox}_{C},S_{L},S_{C},\text{IoU}_{(L,C)},[\text{IoU}_{(L,GD)}]\}$
where ${Bbox}_{L}$ and ${Bbox}_{C}$ are the dimensions of bounding boxes of LiDAR and camera models respectively, represented by their width, height, and center coordinates, normalized by the image’s dimensions. ${S}_{L}$ and ${S}_{C}$ are their confidence scores, and $IoU_{(L,C)}$ is the Intersection over Union of LiDAR and camera detection. When two camera methods are used, information about a second bounding box is provided. For example, $IoU_{(L,GD)}$ is the optionally provided IoU between LiDAR and GroundingDINO.
The entire input vector is then fed to a two-layer multi-layer perceptron (MLP) and trained with a high recall to provide a bias towards avoiding accidentally removing true positive bounding boxes while reducing false positives. This MLP is simple and architecturally identical in all of our experiments. However the activation weights need to be trained for each pair of LiDAR and camera methods since each baseline method has its own distributions of true and false positives and the goal of the MLP is to make specific tradeoffs in these distributions.
In general we desire over production of bounding boxes in the verification method as opposed to under production. Since the detected boxes of the open vocabulary model are not as precise as the specialized LiDAR and camera models, especially around occluded objects, we use a lower IoU threshold to allow more boxes to match.
Easy Hard
Models Modality TP FP TP FP
PVRCNN L 1433 1287 5364 3620
PV+Mono (Ours) LC 1428 393 5280 1084
PV+YOLO (Ours) LC 1430 360 5278 965
PV+GDino (Ours) LC 1430 366 5271 937
PV+Mono+GDino (Ours) LCC 1430 354 5261 906
PV+YOLO+GDino (Ours) LCC 1430 347 5266 887
TED L 1434 864 5425 1848
TED+Mono (Ours) LC 1434 323 5210 599
TED+YOLO (Ours) LC 1434 389 5380 1028
TED+GDino (Ours) LC 1434 352 5359 869
TED+Mono+GDino (Ours) LCC 1434 334 5343 826
TED+YOLO+GDino (Ours) LCC 1434 381 5354 914
TABLE I: Since ours is a late fusion method that generates no new detections, we focus on reducing false positives (FPs) while preserving true positives (TPs). Our method is successfully able to do so, reducing the FPs by an average of 73.6% on PVRCNN and 54.2% on TED.
Figure 5: Our fusion method is able to reduce the false positives significantly on (a) PVRCNN, and (b) TED model detections. Note that this improvement is achieved for all tested camera methods used for verification.
IV-B Implementation
To implement our late-fusion method, we combine two specialized LiDAR detectors (PVRCNN and TED) with various camera models: a specialized model (MonoDETR), a general-purpose model (YOLO-NAS), and an open-vocabulary model (GroundingDINO). The experiments are conducted on the KITTI dataset. We remove dataset images for which ground truth is not available, as well as those which have already been used to train our baseline models. This leaves 3769 images available for our experiments comprising the subset that the dataset providers call validation. We split these images into training and testing subsets. For PVRCNN, these images generate a training set of 12,303 LiDAR bounding box samples and test set of 12,817 bounding boxes. For TED-S, we obtain 9,016 training and 9,406 testing bounding boxes.
The final inputs to our model include details such as LiDAR bounding box dimensions, confidence scores, and IoUs from the camera methods. These inputs pass through a two-layer multi-layer perceptron (MLP) with ReLU activation function and a last sigmoid layer. This output decides whether to keep each LiDAR box or discard the detection. The training loss function compares these outputs with the ground truth labels. Since our focus is on reducing false positives without significantly disturbing the true positive count, we use class weights of 1:10 for negative and positive classes to maintain high recall. We train our models for 50 epochs with a binary cross entropy (BCE) loss and an Adam learning rate of 0.0001.
V Results
We present the results of our methods in two parts: First, we show the impact of our fusion method on reducing false positives generated by LiDAR-only models. Second, we provide an overall comparison of the 2D Average Precision (AP), comparing our fusion method with single-modality models as well as state-of-the-art fusion methods.
V-A False Positives Reduction
Table I shows the performance of LiDAR-only models in terms of true positive and false positive bounding box predictions. Note that both LiDAR models produce a large number of false positives. Our late fusion method is evaluated on each LiDAR model, using each of our three baseline camera detectors, as well as using combinations of camera detectors. Our method consistently and substantially reduces false positives across all combinations. This is made especially clear through the visualization in Figure 5. Our late fusion method achieves an average false positive reduction of 73.6% on PVRCNN and 54.2% on TED. Importantly, the true positive (TP) numbers have only a small reduction, indicating that the significant decrease in false positives does not come at the cost of many missed detections.
2D AP (%)
Models Modality Model Type Easy Moderate Hard
MonoDETR[48] C Specialized 95.12 87.36 79.94
YOLO-NAS[29] C General 75.32 68.69 55.11
GroundingDino[16] C Open-Vocab 84.73 78.68 62.82
PVRCNN[31] L Specialized 98.30 94.44 94.08
Ours (PVRCNN+MonoDETR) LC Late-Fusion 98.88 95.11 92.89
Ours (PVRCNN+YOLO-NAS) LC Late-Fusion 98.88 95.14 92.91
Ours (PVRCNN+GDino) LC Late-Fusion 98.88 95.14 92.92
Ours (PVRCNN+MonoDETR+GDino) LCC Late-Fusion 98.90 95.20 92.96
Ours (PVRCNN+YOLO-NAS+GDino) LCC Late-Fusion 98.92 95.23 92.95
TED-S[41] L Specialized 98.88 96.73 94.96
Ours (TED+MonoDETR) LC Late-Fusion 99.14 93.34 93.35
Ours (TED+YOLO-NAS) LC Late-Fusion 99.15 93.35 95.41
Ours (TED+GDino) LC Late-Fusion 99.11 95.35 95.44
Ours (TED+MonoDETR+GDino) LCC Late-Fusion 99.18 95.46 95.51
Ours (TED+YOLO-NAS+GDino) LCC Late-Fusion 99.13 95.34 95.39
CLOCs [24] LC Late-Fusion 99.33 93.75 92.89
CMKD[9] LC Knowledge Distil. 96.08 92.43 87.11
LogoNet[13] LC Deep Fusion 97.04 89.44 89.20
TED-M[41] LC Deep Fusion 99.48 96.25 93.72
TABLE II: Perfomance on the Average Precision (AP) metric for our method on all difficulty sets on KITTI-val. Our late fusion method consistently outperforms the original specialized LiDAR and camera methods on the easy set, along with moderate on PVRCNN and hard on TED. Our method is also comparable to the SOTA deep fusion methods, despite being simpler and lightweight, without requiring special integrated training.
Figure 6: PRC curves showing the LiDAR method TED before and after adding validation with our method. Especially for the lower confidence predictions needed to produce high recall rates, the LiDAR method produces many false positives. Zoomed-in plots reveal our method’s effectiveness in reducing these false positives, increasing the average precision (AP).
V-B Average Precision
Table II presents the 2D Average Precision (AP) across the Easy, Moderate, and Hard KITTI sets, comparing individual LiDAR models, camera models, and their fusion combinations. Our late fusion method shows improved performance compared to any single-modality models, outperforming PV-RCNN on the Easy and Moderate sets, and TED-S on the Easy and Hard sets. The general and open-vocabulary camera models YOLO-NAS and GroundingDINO, perform as well as the special purpose camera model MonoDETR. These models, not specifically trained on KITTI, achieve comparable performance when fused with LiDAR models, demonstrating the adaptability of our method across different types of detectors.
In comparison with several state-of-the-art (SOTA) fusion methods, our method shows comparable results. Some of the comparison methods show better performance for the Easy and Moderate sets, while our method demonstrates better results on the Hard set. Our late-fusion method can match more complex fusion methods in terms of accuracy while offering better adaptability and lower computational overhead.
To better understand our verification based method, Fig. 6 shows the Precision-Recall (PR) curves comparing TED-S with two of our fusion approaches. The PR curves show that at high recall values our method is able to increase precision by removing false positives.
VI Limitations
Our late fusion method is targeted at reducing false positives, but does not seek to increase true positives. We thus require that the LiDAR (or primary) detector successfully finds most true positives. Similarly, since our approach relies on validating LiDAR detections with camera inputs, it may not perform well with camera models that produce few detections, and is therefore more applicable to methods which can be tuned for a greater number.
VII Conclusion
In this paper, we propose a model-independent late fusion method that combines LiDAR and camera detections to enhance object detection performance for autonomous vehicles. Our approach focuses on reducing false positives by cross-validating LiDAR detections with camera inputs. Experiments on the KITTI dataset demonstrate that our method is exceptional at reducing false positives, and performs competitively with state-of-the-art methods on average precision. Importantly, the method works equally well with several camera detectors, including non-specialized and open-vocabulary models.